OpenStudy (anonymous):

An airplane flew 3.5 hours with a 60 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air.

OpenStudy (anonymous):

d_1 = 3.5(x - 60) d_2 = 2.5(x + 60 d_1 = d_2 3.5(x - 60) = 2.5(x + 60) 3.5x - 210 = 2.5x + 150 x = 360 The speed in still air is 360 mph. These questions often involve the formula s = d/t. You just have to manipulate the formula to change it. s = d/t d = st t = d/s